How can i solve this by using the substitution method?
x=(1/5)y
3x+2y=26

Answer by stanbon(75887)

The 1st equation tell you what "x" is.
Substitute that value of "x" into the 2nd equation to get:
3((1/5)y)+2y=26
(3/5)y+2y=26
(13/5)y=26
(1/5)y=2
y=10
Substitute that back into the 1st equation to find "x".
x=(1/5)y
x=(1/5)(10)
x=2
Solution x=2 and y=10
Cheers,
Stan H.